# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v1

@@ 6 Related Work
Empirical and theoretical evidence shows that tokenization is a central determinant of both the quality and the speed with which Transformers learn. Early byte or character-level systems such as CANINE Clark et al. (2022) and ByT5 Xue et al. (2022) avoid sub-word vocabularies but pay a steep price: sequences become an order of magnitude longer, gradients are noisier, and convergence is markedly slower than for sub-word models. Enlarging the BPE dictionary lets even an unigram model approach near-optimal cross-entropy by approximating word-level tokens, whereas the same model without a tokenizer underfits (Rajaraman et al., 2024). Furthermore, Controlled scaling studies reveal a log–linear pattern: exponentially expanding the input vocabulary leads to an almost linear drop in loss across model sizes Tao et al. (2024). Huang et al. (2025) push this further with “Over-Tokenized Transformers,” showing that a 400400M parameter model with a 12.812.8M token encoder matches the loss of a 11B model baseline without extra compute.
@@
Beyond the gains from larger vocabularies, a growing body of work shows that permitting merges across word boundaries (i.e., disabling whitespace pretokenization) improves language model performance. Liu et al. (2025); Schmidt et al. (2025) turn off white-space pretokenization during BPE tokenizer training once the vocabulary reaches $t$, or at frequency-driven transition points. Taken together, these results indicate that without tokenization, language models stall at character granularity, but with a tokenizer, they converge faster and generalize better.
Transformer language model as a context-aware lossless compressor
Language modeling and lossless compression are effectively two views of the same task: modelling the next token distribution as sharply as possible so that an arithmetic or asymmetric-numeral coder can encode the symbol stream near its entropy rate Shannon (1951); Delétang et al. (2024); Huang et al. (2024). This rationale came from Shannon’s source–coding theorem (Shannon, 1948), which states that the optimal average code-length for text sampled from a distribution $p_{\mathrm{data}}$ is $\mathbb{E}_{x\sim p_{\mathrm{data}}}\!\bigl{[}-\log_{2}p_{\mathrm{data}}(x)\bigr{]}$, the smallest attainable number of bits for losslessly representing $x\sim p_{\mathrm{data}}$. When $p_{\mathrm{data}}$ is known, arithmetic coding (Rissanen, 1976; Pasco, 1977) achieves this bound by converting probabilities into bit-exact codes. In practice, however, $p_{\mathrm{data}}$ is unknown, so we substitute a language model $p_{\mathrm{model}}(x)$ as an approximation. Passing these model-derived probabilities to an arithmetic coder yields efficient compression. Consequently, lowering a language model’s cross-entropy loss is equivalent to building a lossless compressor for the training corpus, with a Transformer acting as a context-sensitive probability estimator that drives the coding process.
@@ 7 Conclusion
8 Limitations
Scaling up model size not only replicates the uncertainty reduction for frequent-word predictions achieved by a larger vocabulary, but also prevents the loss explosion on rare tokens, a pitfall of expanding the vocabulary size that ultimately yields diminishing returns in cross-entropy reduction. However, which scaling dimension – adding depth with extra layers, increasing width by enlarging hidden dimensions, or some other factor – drives this gain, and why, remains an open question. Answering these questions is crucial not only for understanding where the power of larger language models comes from but also for asking whether a carefully designed smaller language model trained from scratch can reproduce the performance of its much larger counterpart.
@@ Appendix A Influence of token-frequency imbalance on unigram and language models
In this section, we provide a detailed explanation of our research question. Natural language exhibits strong contextual dependencies rather than behaving as an i.i.d. process: each token’s probability depends on its preceding context. As a result, the entropy rate, $H_{\infty}\;=\;\lim_{t\to\infty}H\bigl{(}X_{t}\mid X_{<t}\bigr{)}$, which captures the optimal per-token uncertainty in text, is strictly lower than the unigram Shannon entropy $H_{1}\;=\;-\sum_{w}p(w)\,\log p(w)$. Although entropy rate and Shannon entropy coincide for truly i.i.d. data, in natural language, they can differ by several bits per token.
Under a pure unigram model, minimum cross-entropy loss exactly equals Shannon entropy, and modifying the vocabulary size of the tokenizer immediately changes the Shannon entropy. Typically, enlarging the vocabulary segments frequent multi-token patterns into single tokens, driving their individual relative frequency up and reducing Shannon entropy. But if the vocabulary size grows too large, the new entries tend to be rare tokens, which lengthen the tail and can actually increase the Shannon entropy as token-frequency imbalance rises Zouhar et al. [2023]. However, experimental results show that expanding a BPE vocabulary to around $80\mathrm{k}$ lowers the Shannon entropy of unigram models, demonstrating that a more skewed token-frequency distribution is advantageous at practical vocabulary scales Rajaraman et al. [2024]. This pattern extends to n-gram models as well: their conditional Shannon entropy—the lowest possible n-gram cross-entropy—can never exceed the unigram entropy, so lowering the unigram entropy necessarily lowers the conditional entropy.
@@
where $V$ denotes the tokenizer’s vocabulary, $p(x)$ the marginal probability of token $x$, $P(\,\cdot\mid x_{<t})$ the true next-token distribution given the full history $x_{<t}$, and $Q_{\theta}(\,\cdot\mid x_{<t})$ the model’s predicted distribution with parameters $\theta$. When the target label is a one-hot vector, the language model loss can be written as $\mathcal{L}(\theta)=\sum_{x\in V}p(x)\,\bigl{[}-\log Q_{\theta}(x\mid x_{<t})\bigr{]}$ so it is a marginal-frequency-weighted average of the model’s surprisal $\log Q_{\theta}$. By contrast, the Shannon entropy of the unigram model is a marginal-frequency-weighted average of the self-information $\log p(x)$ in the dataset. Even though frequent token logits and embedding norms are higher than rare ones (see §4.1 and Appendix D), the loss $\log Q_{\theta}$ depends not only on the underlying relative token-frequencies but also on training dynamics as well. We therefore cannot derive a closed-form expression to predict how much the loss on frequent tokens shrinks versus how much the rare token losses grow under $\mathcal{L}=p_{\mathrm{freq}}\,L_{\mathrm{freq}}+p_{\mathrm{rare}}\,L_{\mathrm{rare}}$. This masking effect makes it substantially harder to measure the influence of token-frequency imbalance in language models than in the unigram and n-gram models.
@@ Appendix B Coverage of the most frequent N words
In this section, we measure the coverage of the frequent words in Fineweb-edu Penedo et al. [2024] and OpenWebText Gokaslan et al. [2019]. Both dataset exhibit a steep rise in cumulative coverage as we include more high-frequency tokens, but with subtly different baselines and slopes. In figure 6(a), the most frequent $500$ words already cover about 5858% of all tokens, climbing to roughly 7575% once we take the $2,500$ most frequent words. OpenWebText (Figure 6(b)) starts marginally higher—around 5959% at most frequent $500$ words, but follows an almost identical trajectory, reaching about 7676% coverage by the frequent $2,500$ words. This pattern underscores how a relatively small core vocabulary captures the vast majority of running text in both corpora, with only modest gains as we move deeper into the long tail.
@@ Appendix C OpenWebText experiments results
Figure 7: Figure 7(a) reveals that expanding the vocabulary from $24\mathrm{K}$ to $196\mathrm{K}$ steadily reduces the average per-vocabulary loss of high frequency words. Figure 7(b) indicates that the most frequent $2,500$ tokens still contribute roughly 7575% of the total loss, while the rare words losses grow with vocabulary size, similar to 2(b). Figure 7(b) further shows that the global cross-entropy loss falls by about 0.25 nats as the vocabulary grows, demonstrating that the reduction of loss on frequent words outweighs the inflation of rare-token losses.
To verify that reducing frequent-word loss is not a by-product of dataset quality, we repeat the same experiments in section 3.6 on the OpenWebText dataset. Figure 7(a) shows that widening the vocabulary from $24\mathrm{K}$ to $196\mathrm{K}$ in OpenWebText progressively reduces the average loss assigned to high-frequency words. Figure 7(b) indicates that the most frequent $2,500$ tokens still account for about 7575% of the total loss, whereas the loss on infrequent tokens grows with vocabulary size, paralleling the pattern seen in Figure 2(b). Figure 7(b) further demonstrates that the global cross-entropy loss falls from $3.37$ nats at a $24\mathrm{K}$ vocabulary to $3.12$ nats at $196\mathrm{K}$, implying that the reduction in loss on frequent words more than offsets the increase in rare-token loss, regardless of dataset quality or type.
Appendix D Frequent and rare token norm in output embedding
In this section, we explain why frequent tokens acquire larger output-embedding norms and logits by deriving and analysing the gradients of the output embedding. Cross-entropy loss with a vocabulary size $|V|$ and hidden dimension size $d_{\text{model}}$, using a one-hot target $t\in\mathbb{R}^{|V|}$, the logit vector is $\ell=hE_{\text{out}}^{\top}$ where $h\in\mathbb{R}^{d_{\text{model}}}$ is the final hidden state and $E_{\text{out}}\in\mathbb{R}^{|V|\times d_{\text{model}}}=[u_{1},\dots,u_{|V|}]$ is the output-embedding matrix. When input and output embeddings are untied, the gradient with respect to each row of $E_{\text{out}}$ decomposes into
$\frac{\partial\mathcal{L}}{\partial E_{\text{out}_{t}}}=(p_{t}-1)\,h,\qquad\frac{\partial\mathcal{L}}{\partial E_{\text{out}_{j}}}=p_{j}\,h\;\;(j\neq t),$ (6)
where $p_{t}=\mathrm{softmax}(\ell)_{t}$ Bis et al. [2021], Mircea et al. [2024]. Because $p_{t}\ll 1$ at the start of training, $\bigl{\|}\partial\mathcal{L}/\partial E_{\text{out}_{t}}\bigr{\|}_{2}\approx\|h\|_{2}$ while each non-target row scales only with $p_{j}\|h\|_{2}$ ($p_{j}\!<\!1/|V|$). Thus, every time token $t$ appears, its embedding is pulled almost $\|h\|_{2}$ units along +h+h, whereas each competing row is nudged by a factor of $p_{j}\!\ll\!1$. As tokens recur in the training data, their token embeddings accumulate gradient updates roughly proportional to their counts, so $\|E_{\text{out}_{t}}\|_{2}$ grows in line with token-frequency. Since each logit factorizes as $\ell_{t}=\|h\|_{2}\,\|E_{\text{out}_{t}}\|_{2}\cos\theta_{t}$ and frequent tokens not only acquire the largest norms but also align closely with final hidden-state directions $\cos\theta_{t}\approx 1$, they end up with disproportionately large logits whereas rare tokens suffer both smaller norms and larger angular deviations Bis et al. [2021], Land and Bartolo [2024]. Although $\ell_{2}$ weight decay can slow this norm inflation, they merely dampen the effect rather than remove the underlying frequency norm logit correlation. This frequency-proportional amplification explains the empirical observation that output-embedding norms and softmax logits scale with token-frequency in standard language models.
Appendix E OLMo-2 result
@@
Figure 8: Figure 8(a) illustrates that larger $7\mathrm{B}$ model reduces average per-vocabulary loss from $3.39$ nats to $3.07$ nats while slightly increasing the proportion of loss covered by frequent words from 7171% to $72.5\%$. Figure 8(b) further demonstrates that Scaling from $1\mathrm{B}$ to $7\mathrm{B}$ reduces the overall cross-entropy from $2.35$ nats to $2.12$ nats, confirming the same pattern persists in contemporary large-scale language models.
@@
To identify whether the reduction in frequent words loss with increasing model size holds for contemporary large language models, we perform analogous experiments using the OLMo-2 series OLMo et al. [2025]. Figure 8(a) and 8(b) indicate that the average per vocabulary loss falls from $3.39$ nats in the 11B parameter model to $3.07$ nats in the 77B variant while slightly increasing the proportion of loss covered by $2,500$ frequent words from 7171% to $72.5\%$. Figure 8(b) further shows that global cross-entropy loss declines from $2.35$ nats for OLMo-2 $1\mathrm{B}$ to $2.12$ nats for OLMo-2 77B. Notably, OLMo-2 employs a much larger vocabulary (cl100K OpenAI [2023]) than Pythia (50304 tokens Black et al. [2022]) and trains on a larger corpus OLMo et al. [2025], which helps drive down the average loss on high-frequency words. These results confirm that the same trend holds for modern large-scale language models as well.

### v2

@@ 6 Related Work
Empirical and theoretical evidence shows that tokenization is a central determinant of both the quality and the speed with which Transformers learn. Early byte or character-level systems such as CANINE [9] and ByT5 [60] avoid sub‑word vocabularies but pay a steep price: sequences become an order of magnitude longer, gradients are noisier, and convergence is markedly slower than for sub‑word models. Enlarging the BPE dictionary lets even an unigram model approach near-optimal cross-entropy by approximating word-level tokens, whereas the same model without a tokenizer underfits [44]. Furthermore, Controlled scaling studies reveal a log–linear pattern: exponentially expanding the input vocabulary leads to an almost linear drop in loss across model sizes [55]. Huang et al. [20] push this further with “Over‑Tokenized Transformers,” showing that a 400400M parameter model with a 12.812.8M token encoder matches the loss of a 11B model baseline without extra compute.
@@
Beyond the gains from larger vocabularies, a growing body of work shows that permitting merges across word boundaries (i.e., disabling whitespace pretokenization) improves language model performance. Liu et al. [32], Schmidt et al. [49] turn off white-space pretokenization during BPE tokenizer training once the vocabulary reaches $t$, or at frequency-driven transition points. Taken together, these results indicate that without tokenization, language models stall at character granularity, but with a tokenizer, they converge faster and generalize better.
Large language model performs a two stage lossless compression
Large language models perform a two stage lossless compression where tokenizer acts as a pre-compressor and transformer models the tokenized data [11, 30]. Lester et al. [30] further explore the role of tokenizer as a pre-compressor by replacing it with small transformer model with arithmetic coding, and models trained on neurally compressed text underperforms subword tokenizers but beat byte-level LMs. Delétang et al. [11] and Valmeekam et al. [57] has been proposed that transformer language models can be harnessed for compression via arithmetic coding and Heurtel-Depeiges et al. [19] extends this line by probing the possibility of transformers as a universal compressor across text, images, and audio. However, most literatures mixuse the concept of generalization and compression where generalization focuses on predicting unseen data from different datasets by exploiting shared statistical patterns while compression focuses on encoding a fixed dataset in as few bits as possible by modeling its distribution and removing redundancy. This work investigates the role and impact of tokenizer on generalization through the lens of data compression and information theory.
BPE tokenizer with inductive biases
BPE tokenizers typically adopt pre-tokenization where regular expressions split text into chunks, sometimes called pre-tokens. Recent work shows that removing pre-tokenization degrades downstream performance across diverse tokenizers [10, 47]. Pre-tokenization rules act as an inductive bias grounded in morphology and grammar, helping the tokenizer capture ubiquitous statistical patterns in natural text, which is crucial for transformer language models to generalize unseen data.
@@ 7 Conclusion
Acknowledgement
This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-II220184, 2022-0-00184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics).
@@ Appendix A Influence of token-frequency imbalance on unigram and language models
In this section, we provide a detailed explanation of our research question. Natural language exhibits strong contextual dependencies rather than behaving as an i.i.d. process: each token’s probability depends on its preceding context. As a result, the entropy rate, $H_{\infty}\;=\;\lim_{t\to\infty}H\bigl(X_{t}\mid X_{<t}\bigr)$, which captures the optimal per‐token uncertainty in text, is strictly lower than the unigram Shannon entropy $H_{1}\;=\;-\sum_{w}p(w)\,\log p(w)$. Although entropy rate and Shannon entropy coincide for truly i.i.d. data, in natural language, they can differ by several bits per token.
Under a pure unigram model, minimum cross‐entropy loss exactly equals Shannon entropy, and modifying the vocabulary size of the tokenizer immediately changes the Shannon entropy. Typically, enlarging the vocabulary segments frequent multi-token patterns into single tokens, driving their individual relative frequency up and reducing Shannon entropy. But if the vocabulary size grows too large, the new entries tend to be rare tokens, which lengthen the tail and can actually increase the Shannon entropy as token-frequency imbalance rises [64]. However, experimental results show that expanding a BPE vocabulary to around $80\mathrm{K}$ lowers the Shannon entropy of unigram models, demonstrating that a more skewed token-frequency distribution is advantageous at practical vocabulary scales [44]. This pattern extends to n-gram models as well: their conditional Shannon entropy—the lowest possible n-gram cross-entropy—can never exceed the unigram entropy, so lowering the unigram entropy necessarily lowers the conditional entropy.
@@
where $V$ denotes the tokenizer’s vocabulary, $p(x)$ the marginal probability of token $x$, $P(\,\cdot\mid x_{<t})$ the true next-token distribution given the full history $x_{<t}$, and $Q_{\theta}(\,\cdot\mid x_{<t})$ the model’s predicted distribution with parameters $\theta$. When the target label is a one-hot vector, the language model loss can be written as $\mathcal{L}(\theta)=\sum_{x\in V}p(x)\,\bigl[-\log Q_{\theta}(x\mid x_{<t})\bigr]$ so it is a marginal‐frequency‐weighted average of the model’s surprisal $\log Q_{\theta}$. By contrast, the Shannon entropy of the unigram model is a marginal‐frequency‐weighted average of the self-information $\log p(x)$ in the dataset. Even though frequent token logits and embedding norms are higher than rare ones (see Appendix LABEL:apdx:embdnorm and LABEL:apdx:unembd_grad), the loss $\log Q_{\theta}$ depends not only on the underlying relative token-frequencies but also on training dynamics as well. We therefore cannot derive a closed-form expression to predict how much the loss on frequent tokens shrinks versus how much the rare token losses grow under $\mathcal{L}=p_{\mathrm{freq}}\,L_{\mathrm{freq}}+p_{\mathrm{rare}}\,L_{\mathrm{rare}}$. This masking effect makes it substantially harder to measure the influence of token‐frequency imbalance in language models than in the unigram and n-gram models.
@@ Appendix B Coverage of the most frequent N words
In this section, we measure the coverage of the frequent words in Fineweb-edu [40] and OpenWebText [13]. Both dataset exhibit a steep rise in cumulative coverage as we include more high-frequency tokens, but with subtly different baselines and slopes. In figure 6(a), the most frequent $500$ words already cover about 5858% of all tokens, climbing to roughly 7575% once we take the $2,500$ most frequent words. OpenWebText (Figure 6(b)) starts marginally higher—around 5959% at most frequent $500$ words, but follows an almost identical trajectory, reaching about 7676% coverage by the frequent $2,500$ words. This pattern underscores how a relatively small core vocabulary captures the vast majority of running text in both corpora, with only modest gains as we move deeper into the long tail.
@@ Appendix C OpenWebText experiments results
Figure 7: Figure 7(a) reveals that expanding the vocabulary from $24\mathrm{K}$ to $196\mathrm{K}$ steadily reduces the average per-vocabulary loss of high frequency words. Figure 7(b) indicates that the most frequent $2,500$ tokens still contribute roughly $75$ of the total loss, while the rare words losses grow with vocabulary size, similar to 2(b). Figure 7(b) further shows that the global cross-entropy loss falls by about $0.25$ nats as the vocabulary grows, demonstrating that the reduction of loss on frequent words outweighs the inflation of rare-token losses.
To verify that reducing frequent-word loss is not a by-product of dataset quality, we repeat the same experiments in section 3.6 on the OpenWebText dataset. Figure 7(a) shows that widening the vocabulary from $24\mathrm{K}$ to $196\mathrm{K}$ in OpenWebText progressively reduces the average loss assigned to high-frequency words. Figure 7(b) indicates that the most frequent $2,500$ tokens still account for about $75$ of the total loss, whereas the loss on infrequent tokens grows with vocabulary size, paralleling the pattern seen in Figure 2(b). Figure 7(b) further demonstrates that the global cross-entropy loss falls from $3.37$ nats at a $24\mathrm{K}$ vocabulary to $3.12$ nats at $196\mathrm{K}$, implying that the reduction in loss on frequent words more than offsets the increase in rare-token loss, regardless of dataset quality or type.
Appendix D OLMo-2 result
@@
Figure 8: Figure 8(a) illustrates that larger $7\mathrm{B}$ model reduces average per-vocabulary loss from $3.39$ nats to $3.07$ nats while slightly increasing the proportion of loss covered by frequent words from $71\%$ to $72.5\%$. Figure 8(b) further demonstrates that Scaling from $1\mathrm{B}$ to $7\mathrm{B}$ reduces the overall cross-entropy from $2.35$ nats to $2.12$ nats, confirming the same pattern persists in contemporary large-scale language models.
@@
To identify whether the reduction in frequent words loss with increasing model size holds for contemporary large language models, we perform analogous experiments using the OLMo-2 series [38]. Figure 8(a) and 8(b) indicate that the average per vocabulary loss falls from $3.39$ nats in the 11B parameter model to $3.07$ nats in the 77B variant while slightly increasing the proportion of loss covered by $2,500$ frequent words from $71\%$ to $72.5\%$. Figure 8(b) further shows that global cross-entropy loss declines from $2.35$ nats for OLMo-2 $1\mathrm{B}$ to $2.12$ nats for OLMo-2 77B. Notably, OLMo-2 employs a much larger vocabulary (cl100K [39]) than Pythia (50304 tokens [6]) and trains on a larger corpus [38], which helps drive down the average loss on high-frequency words. These results confirm that the same trend holds for modern large-scale language models as well.
Appendix E Are the gains mainly driven by parameter increases from vocab expansion?
To probe whether gains come from parameter growth with larger vocabularies, we run a controlled study comparing a $24\mathrm{K}$ vocabulary model ($122\mathrm{M}$ parameters) to a near equal size $49\mathrm{K}$ model ($124\mathrm{M}$ parameters) to isolate embedding capacity effects. We reduced the hidden dimension from $768$ to $648$ and the feed‑forward dimension from $2048$ to $1728$, while maintaining identical model depth which known to have a dominant impact on performance [56, 41]. As shown in the table 3, the $49\mathrm{K}$ model incurs higher cross‑entropy loss and delivers inferior results on downstream benchmarks (ARC‑E, HellaSWAG, PIQA, and SCIQ) compared to its $24\mathrm{K}$ counterpart, indicating that simply expanding the embedding layer contributes far less to model expressivity than is often assumed.
Table 3: A $24\mathrm{K}$ vs. size-matched $49\mathrm{K}$ indicates embedding expansion yields limited gains.
Vocab (Params) CE Loss Downstream (%)
24K (122M) 3.179 54.73
49K (124M) 3.563 50.68
49K (161M) 3.171 55.44
Appendix F Do the benefits of larger vocabularies persist with tied-embedding models?
Several studies show that tying input and output embeddings can change pre-training dynamics [22, 29]. We therefore repeat our study with tied embeddings and report average frequent-word loss and cross-entropy on FineWeb-Edu in the table 4. The tied embedding variant shows broadly similar average frequent word loss and ce-loss to the untied baseline.
Table 4: Average frequent-word loss and cross-entropy of tied and untied embedding models
24K 49K 98K 196K
Tied Avg frequent word loss 4.067 4.029 3.991 3.968
Cross entropy 3.177 3.160 3.146 3.134
Untied Avg frequent word loss 4.074 4.031 4.001 3.977
Cross entropy 3.179 3.163 3.147 3.136
Appendix G To what extent does a larger model decrease frequent and rare word loss?
We conducted experiments to quantify scaling effects on frequent and rare words loss and demonstrate that scaling model capacity disproportionately benefits the prediction of high‑frequency tokens relative to low‑frequency tokens. Since BPE vocabulary IDs are not token frequency ordered, we cannot directly distinguish common and rare tokens in the Pythia and OLMo‑2 training corpora. To proxy frequency, we utilize the Google Web Trillion Word Corpus frequency list, which records counts for $333,333$ unique words appearing at least $10,000$ times in a $1.0249\times 10^{12}$ words. We classify the top $2,500$ entries as frequent words and the bottom $10,000$ as rare words. As shown in the table 5, although cross‑entropy loss declines for both groups with increasing model size, the reduction for frequent tokens is markedly greater. We therefore conclude that the dominant effect of model scaling is the lowering of frequent‑token loss, which in turn drives the overall cross‑entropy reduction.
Table 5: Frequent, rare word loss and cross-entropy loss of Pythia and OLMo-2
Model Params Frequent word loss Rare word loss CE-loss
Pythia 1B 3.01 4.99 2.51
7B 2.55 4.93 2.26
OLMo-2 1B 2.60 4.27 2.22
7B 2.20 4.19 2.00
Appendix H Learning rate exploration
We use a fixed learning rate of $6\times 10^{-4}$ for an $85\mathrm{M}$ non-embedding–parameter language model, mirroring the learning rate of GPT-3 with an equivalent non-embedding model size [7]. we further conducted a learning‑rate exploration with four additional learning rates: $2.4\times 10^{-3}$, $1.2\times 10^{-3}$, $1.5\times 10^{-4}$, $7.5\times 10^{-5}$. Table 6 and 7 report the resulting cross‑entropy loss and average word loss of frequent 2500 words in fineweb-edu across different vocabulary sizes and learning rates, using the same validation dataset described in Section 3.1.
Table 6: Cross-entropy loss across five learning rates.
Cross-Entropy Loss 2.4e-3 1.2e-3 6.0e-4 1.5e-4 7.5e-5
24K 3.156 3.141 3.179 3.484 3.779
49K 3.135 3.114 3.163 3.461 3.751
98K 3.122 3.099 3.147 3.443 3.736
196K 3.111 3.086 3.136 3.429 3.725
Table 7: Average frequent-word loss across five learning rates.
Frequent Word Loss 2.4e-3 1.2e-3 6.0e-4 1.5e-4 7.5e-5
24K 4.033 4.012 4.074 4.431 4.783
49K 3.990 3.970 4.031 4.399 4.749
98K 3.963 3.939 4.001 4.373 4.725
196K 3.942 3.916 3.977 4.352 4.703
Appendix I Detailed Experimental Setting
In this section, we provide detailed configurations of pretraining to reproduce our results. The training setup (Table 8) uses a global batch size of 256, weight decay 0.1, and sequence length 2048. Optimization is Adam with a cosine learning-rate schedule, a $700$-step warmup, and a weight-initialization scale of 0.02. The model setup (Table 9) covers two different model size: an 85M model with 12 layers and 12 heads $(d_{\text{model}}=768,\;d_{\text{ffn}}=2048,\;d_{\text{head}}=64)$ and a 450M model with 21 layers and 21 heads $(d_{\text{model}}=1344,\;d_{\text{ffn}}=3548,\;d_{\text{head}}=21)$. Together, these tables specify the standardized training hyperparameters and the core architectural dimensions for both scales.
Table 8: Training configurations. LR Schedule denotes learning-rate schedule.
Global Batch Size Weight Decay Sequence Length Optimizer
256 0.1 2048 AdamW
LR Schedule Warmup Weight Init.
Cosine 700 steps 0.02
Table 9: Model configurations.
Size $n_{\text{layers}}$ $n_{\text{heads}}$ $d_{\text{model}}$ $d_{\text{ffn}}$ $d_{\text{head}}$
85M 12 12 768 2048 64
450M 21 21 1344 3548 21